4. A substance used in biological and medical research is shipped by airfreight to users in cartons of 1,000 ampules. The data below, involving 10 shipments, were collected on the number of times the carton was transferred from one aircraft to another over the shipment route (X) and the number of ampules found to be broken upon arrival (Y). i 12345 67 89 10 X 1 02 03 1 0 2 0 Y 169 17 12 2213 8 15 19 11 (a) Plot the data. (b) Obtain the estimated simple linear regression model. (c) Plot the fitted line on the scatterplot with the data (d) Does a linear regression function appear to give a good fit here? (e) Obtain a point estimate of the expected number of broken ampules when one transfer is made. (f) Obtain the residuals. Do they sum to 0?
Answer #1

We use Excel to solve this question-

Book1 Excel (Product Activation Failed) HOME INSERT PAGE LAYOLT FORMULAS DATA REVIEW EW Sign i Copy or Format Painter Hill B l u .凼. 2. . -_- 4e47 F Merge & Center . s-% , conditional tormatas Neutral Check Cell Insert Delete Format ClestFilter-Selert Sort Hind B rmatting Tahlr Edtng 16 SUMMARY OUTPUT Scatter Plot 17 12 Regressian Statisties Multiple R R Square Adiusted RS U.89 Standard Err1.483 0.95 0.90 13 15 10 19 ANDVA 12 13 ignificonceF 72.727 2.749 05 MS Rcgression 160 17.6 Tutal 17 19 19 CoefficientsStandord Errort Stot p-value Lower 95% 8.b/0 Upper 95% 15,377 8.528 Lower 95.0% Upper 95.0% 11.V30 5.082 Intercept 10.2 0,653 O.UDO 0.000 11.130 .082 8.b 0 2.918 Variable 1 0.469 21 27 RESIDUAL OLTPLT 21 Ohserutian Predicted Y Rsidsas 10.2 18.2 10.2 2.2 20 29 1.2 1.8 0.2 Sheel| | Sheet2 :52 PM O Type here to searchBook1 Excel (Product Activation Failed) HOME INSERT PAGE LAYOLT FORMULAS DATA REVIEW EW cut Copy . Sign i 2 conditional tormatas Neutral Check Cell Insert Delete Format B l u .凼. 2. . -_- 4e47 F Merge & Center . s-% , Sort Hind B Filter. Selert or Format Painter Cleat . rmatting Tahlr Edtng 12 Regressiuni 160 160 72.727 2.749E-05 17.6 2.2 15 16 otal 95.0% Upper nefirients Standand Frmar 0.663 0.169 Srat i ower 95% inner Uppr 9536 Inwr「 11.730 5.082 18 15.377 0.000 8.670 8.670 11.730 X Variable 1 B.528 0.000 5.082 Scatter Plot RESIDUAL OUTPUT Observotion Preaicted Resacuos 1.8 1.2 14.2 10.2 -0.2 1.2 2.2 14.2 10.2 10 10.2 0.8 2.5 Sheel| | Sheet2 ョ1 +101% - :56 PM O Type here to search

a) Please see attachments
b) The estimated simple linear regression model is,
   y^ = 10.2 + 4 * x
c) Please see attachments
d) The coefficient of determination, R square = 0.90 indicate that regression function appear to good fit here.
e) at x = 1, the expected number of broken ampules, y^ is given by,
   y^ = 10.2 + 4 * x
   y^ = 10.2 + 4 * 1
   y^ = 14.2
f) Please see attachments, Yes residuals sum to 0.